OpenStudy (anonymous):

First derivative of Arc sin x/a + sq rt of a^2 - x^2 /x

OpenStudy (anonymous):

\[Arc \sin \frac{ x }{ a } + \frac{ \sqrt{a^2 - x^2} }{ x }\]